# Multi-type branching processes with immigration generated by point processes

### v1

@@ 3.1 Single-type processes
The underlying branching process without immigration $Z_{\times}$ will be a Markov branching process in continuous time with starting state $Z_{\times}(0)=1$. Therefore we can construct the process with immigration $Z$ as follows:
  • At times $T_{1}<T_{2}<\dots$, new particles (immigrants) $I_{1},I_{2},\dots$ arrive with $I_{k}\sim I$ iid. We assume that $\mathbb{E}\left[I^{2}\right]$ is finite.
  • Each particle evolves independently of the others and at $Exp(1/\mu)$ time dies and produces $\nu$ new particles. It is then known, see [2, p.102], that
@@
    where $G_{Z_{\times}}(t,s):=G_{Z_{\times}(t)}(s)$ and $f_{\nu}(s)=(G_{\nu}(s)-s)/\mu$. Also, $M_{\times}(t):=\mathbb{E}\left[Z_{\times}(t)\right]=e^{\rho t}$ with $\rho:=\mu(\mathbb{E}\left[\nu\right]-1)$. The process is called subcritical, critical or supercritical according to whether $\mathbb{E}\left[\nu\right]$ is, respectively, less, equal, or larger than 1.
@@
3.2 Multi-type processes
Consider now a generalisation of the process presented above which is composed of $d$ types of particles. We note that we will use bold symbols, $\boldsymbol{\nu}$ for a vector, and, unless stated otherwise explicitly, $\nu_{i}$ will denote the value of its $i$th coordinate. We define the generating function in this case as
$G_{\boldsymbol{\nu}}(\boldsymbol{s}):=\sum_{\boldsymbol{n}\in\mathbb{N}^{d}}\mathbb{P}\left(\boldsymbol{\nu}=\boldsymbol{n}\right)\prod_{i=1}^{d}s_{i}^{n_{i}}.$
The multi-type branching process $\boldsymbol{Z}$ with $\boldsymbol{Z}(0)=\boldsymbol{I}^{(0)}\sim\boldsymbol{I}$, is described by:
  • At times $0=T_{0}<T_{1}<T_{2}<\dots$, iid immigrants $\boldsymbol{I}^{(0)},\boldsymbol{I}^{(1)},\boldsymbol{I}^{(2)},\dots\sim\boldsymbol{I}$ join the population described by the branching process $\boldsymbol{Z}.$ We assume that $\mathbb{E}\left[I_{i}^{2}\right]$ are finite for all $i$.
  • Each particle of type $i$ lives $Exp(\mu_{i})$ time after which it dies and produces new ones with distribution $\boldsymbol{\nu}_{i}$ of their counts, that is to say $({\nu}_{i})_{j}$ of type $j$.
@@ 3.2 Multi-type processes
  1. (i)
    the immigrants and progeny, ${I_{i}}$ and ${(\nu_{i})}_{j}$ have finite second moments;
  2. (ii)
    the matrix $M:={(\nu_{i})}_{j}$ is such that for some $p\in\mathbb{N}$, $M^{p}$ has strictly positive terms;
  3. (iii)
    and ${(\nu_{i})}_{j}$ have non-lattice distributions.
Similarly to the first case,
@@
where ${Z}^{(i)}_{\times}$ are iid processes with law $\boldsymbol{Z_{\times}}$, that is, without immigration and started with $\boldsymbol{I^{(i)}}$. This gives the generating function of the process, with $G_{\boldsymbol{Z}}(t,\boldsymbol{s}):=G_{\boldsymbol{Z}(t)}(\boldsymbol{s})$ and similarly for $G_{\boldsymbol{Z}_{\times}}(t,\boldsymbol{s}$),
@@
Translating the last in the language of the Laplace transform of the random measure $\Phi$ governing the process $T_{i}$, substituting in (2.1), we get that
$G_{\boldsymbol{Z}}(t,\boldsymbol{s})=\mathcal{L}_{\Phi}(f_{t}),\quad\text{with}\quad f_{t}(x):=-\ln\left(G_{\boldsymbol{Z}_{\times}}\left(t-x,\boldsymbol{s}\right)\right)\mathbbm{1}_{\left\{x\leq t\right\}}.$
@@
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\mathbb{E}\left[\exp(\langle\boldsymbol{Z}(t),\boldsymbol{s}\rangle)\right],$
[…]
@@ 4.1 Single-type processes
A standard argument using martingale theory shows that there is a random variable $W$ such that for the process without immigration $Z_{\times}$, there exists a real-valued random variable $W_{\times}$ such that
@@
In the supercritical case, it is natural to characterise some of the properties of $W_{\times}$, e.g., the existence of a density. In the sub- and critical cases, this $W_{\times}$ is a.s. equal to 00 and a common question is to analyse the so-called Yaglom limit as $t\to\infty$ of quantities such as $\mathbb{P}(Z(t)>f(t)|Z(t)>0)$ for some $f$, see [2, Chapter III, Theorems 2-4]. Similar results are also available for the case of age-dependent processes, see [2, Chapter IV]
@@
    In the case of an immigration spanned by a $(\Lambda,k)$-DPP process, $\mathbb{E}\left[s^{Z(t)}\right]$ is equal to
    $1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\prod_{i=1}^{n}\left(1-G_{Z_{\times}}(t-x_{i},s)\right)\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n}).$ (4.3)
@@
    In the case of an immigration spanned by a Cox process with a directing measure $\eta$,
    $\mathbb{E}\left[s^{Z(t)}\right]={\mathbb{E}_{\eta}\left[\exp\left(-\int_{(0,t]}\left(1-G_{Z_{\times}}(t-x,s)\right)\Lambda(\mathrm{d}x)\right)\right]}.$
@@
    In the case of an immigration spanned by an $\alpha$-FPP,
    $\mathbb{E}\left[s^{Z(t)}\right]={\mathbb{E}_{\eta}\left[\exp\left(-\int_{(0,Y_{\alpha}(t)-]}\left(1-G_{Z_{\times}}\left(t-Y^{-1}_{\alpha}(t),s)\right)\right)\Lambda(\mathrm{d}t)\right)\right]}.$
@@
For the first item, we substitute the Laplace transform of DPPs (2.2) in (3.2). For the second, we use that Cox processes are Poisson ones with a random intensity. For the third item, we use the representation of FPPs as time-changed Poisson processes from (2.3). The last gives that if $\Phi$ is the Laplace transform of an $\alpha$-FPP,
$\mathcal{L}_{\Phi}\left(f\right)=\mathbb{E}\left[\exp\left(-\int\left(1-e^{-f\left(Y_{\alpha}^{-1}(t)\right)}\right)\right)\Lambda(\mathrm{d}t)\right].$
The object $Y_{\alpha}^{-1}(t)$ is the inverse stable subordinator and its analysis is often crucial in understanding the fractional Poisson process after the work [27] showed its relevance. Recently, the asymptotics of inverse subordinators were well understood under very mild assumptions by Ascione et al. in [1]. Substituting in (3.2), we obtain the third item of the theorem. ∎
@@
4.2 Multi-type processes
[…]
$A_{ij}:=\mu_{i}^{-1}\left(\mathbb{E}\left[\left(\nu_{i}\right)_{j}\right]-\mathbbm{1}_{\left\{i=j\right\}}\right).$ (4.4)
We will suppose that $e^{A}$ is irreducible and let $\rho$ be its largest eigenvalue. The process is called sub-, super- or just critical according to $\rho$, respectively, less, larger or equal than 0. Let $\boldsymbol{u}$ and $\boldsymbol{v}$ be row vectors such that
$A\boldsymbol{u}^{t}=\rho\boldsymbol{u}^{t},\quad\boldsymbol{v}A=\rho\boldsymbol{v},\quad\sum u_{i}v_{i}=1,\quad\text{and}\quad\sum_{i}u_{i}=\\ 1,$
with $t$ denoting the transpose. Then, from [2, Theorem 1, p.205] or rewritten in the form [35, Lemma 1], there exists a real-valued random variable $W_{\times}$ such that
@@ 4.2 Multi-type processes
$G_{\boldsymbol{Z}_{\times}}(s)=G_{\boldsymbol{I}}\left(G_{\boldsymbol{Z}^{(1)}_{\times}}(\boldsymbol{s}),\dots,G_{\boldsymbol{Z}^{(d)}_{\times}}(\boldsymbol{s})\right),$
@@
The multidimensional analogue of Theorem 4.1, which is obtained by substituting $s$ and $Z$ with, respectively, $\boldsymbol{s}$ and $\boldsymbol{Z}$, holds.
@@
As noted previously, the case of immigration generated by a Poisson random measure is analysed by Yanev et al. in [17, 42, 43] and independently by Rabehasaina and Woo in [35]. For immigration times that form a renewal process, see [41]. The latter may be applied in the case of an immigration spanned by a FPP after checking that the used arguments can be modified to include the case of an infinite expectation of the inter-arrival times.
As an illustrative example of how ideas from the case of Poisson immigration can be modified to encompass the case of immigration spanned by a DPP, we generalise [35, Theorems 3, 4, and 5] (since the Poisson point process is a DPP), which are also available in [17, 42, 43].
@@ 4.3 Moments of the process
In the case of immigration spanned by a $(\Lambda,K)$-DPP, and denoting the $i$th coordinate with an index,
  1. (i)
@@
This follows by differentiating the Laplace transform of the process, derived as the multidimensional version of (4.3). We provide the complete calculation in Section 5.1.
@@
In the case of immigration spanned by a $(\Lambda,K)$-DPP $\Phi$ with kernel $K$ such that $\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)<\infty$, there exists an $\mathbb{R}^{d}$-valued random variable $\boldsymbol{W}$ such that
[…]
@@
$\boldsymbol{W}\stackrel{{\scriptstyle d}}{{=}}\sum_{i}\boldsymbol{v}\cdot W^{(i)}_{\times}e^{-\rho T_{i}}$
with $W^{(i)}_{\times}$ iid copies of $W_{\times}$ and $T_{i}$ being the atoms of $\Phi$.
@@
where $\boldsymbol{x}+t:=(x_{1}+t,\dots,x_{n}+t)$; see [3, Chapter 5.6]. This property holds, for example, if the kernel is shift-invariant, $K(x,y)=K(x+t,y+t)=:C(x-y)$, as in the case of the Poisson kernel, $K(x,y)=\mathbbm{1}_{\left\{x=y\right\}}$, or the Ginibre (Gaussian) kernel: $K(x,y)=\exp\left(-(x-y)^{2}/2\right)/\pi$. A result by Lavancier et al. [23, Theorem 1] states that $C$ can be any continuous correlation function with eigenvalues in its spectral representation bounded by 1, with examples provided in their work.
@@
Assume that immigration is spanned by a stationary $(\Lambda,K)$-DPP with $\Lambda(\mathrm{d}x)=\lambda(x)\mathrm{d}x$, and that $\lambda(x)\sim\lambda_{\infty}e^{\delta x}$.
@@
    For all $\rho$, if $\delta>\max\{\rho,0\}$, then
    $\frac{{Z_{i}}(t)}{e^{\delta t}}\xrightarrow[t\to\infty]{L^{2}}A_{i}:=K_{\ast}\lambda_{\infty}\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle v_{i}\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x,$
@@
    If the process is supercritical, that is, $\rho>0$, and $\delta=\rho$, then
@@
    Assume that the process critical, i.e. $\rho=0$, $\lambda$ is bounded, $\delta=\rho$, and the covariance function $K(x,0)$ of the DPP tends to 00 as $x\to\infty$. Then
@@
    where $Y\sim Gamma(\lambda_{\infty}\beta,Q)$ with
    $Q:=\frac{1}{2}\sum_{i,j,k=1}^{d}\left.\frac{\partial^{2}G_{\boldsymbol{\nu_{i}}}}{\partial x_{j}\partial x_{k}}\right|_{\boldsymbol{x}=\boldsymbol{1}}v_{i}u_{j}u_{k},\qquad\beta:=\frac{\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle}{Q},$
@@
The result of item 3 in the last theorem in the case of Poisson immigration is established in [17, Theorem 8] and, under the stronger assumption that all moments of $I_{i}$ exist, in [35, Theorem 3]. After a close inspection, we have observed that the two in fact do not match. This is due to two errata: in the proof of [17, Theorem 8], there is a missed $r^{\ast}$ on the second line of page 221. The error in [35, Theorem 3] is due to an incorrect application to the result of Weiner [47]. In his work, the eigenvectors $\boldsymbol{u^{W}}$ and $\boldsymbol{v^{W}}$ are defined as, respectively, the right and left eigenvector of the matrix composed of $\mathbb{E}\left[(\nu_{i})_{j}\right]-\mathbbm{1}_{\left\{i=j\right\}}$, and not as the eigenvectors of the generator matrix $A$, see (4.4). The normalisation is $\langle\boldsymbol{u^{W}},\boldsymbol{1}\rangle=1$ and $\langle\boldsymbol{u^{W}},\boldsymbol{v^{W}}\rangle=1$. It is then direct that,
$u_{i}^{W}=u_{i},\quad v_{i}^{W}=\frac{1}{\sum_{i}(u_{i}v_{i}/\mu_{i})}\frac{v_{i}}{\mu_{i}},\quad\text{and}\quad\sum_{i}\mu_{i}u_{i}^{W}v_{i}^{W}=\frac{1}{\sum_{i}(u_{i}v_{i}/\mu_{i})}.$ (4.8)

### v2

@@ 3.1 Single-type processes
The underlying branching process without immigration $Z_{\times}$ is a Markov branching process in continuous time with starting state $Z_{\times}(0)=I$, where $I$ will be the law of the size of a single immigrant group. Therefore we can construct the process with immigration $Z$ as follows:
  • At times $T_{1}<T_{2}<\dots$, new particles (immigrants) $I_{1},I_{2},\dots$ arrive with $I_{k}\sim I$ iid. We assume that $\mathbb{E}\left[I\right]$ is finite.
  • Each particle evolves independently of the others and at $Exp(1/\mu)$ time dies and produces $\nu$ new particles. It is then known, see [2, (5) on p.106], that
@@
    where $G_{Z_{\times}}(t,s):=G_{Z_{\times}(t)}(s)$ and $f_{\nu}(s)=(G_{\nu}(s)-s)/\mu$. Also, $M_{\times}(t):=\mathbb{E}\left[Z_{\times}(t)\right]=e^{\rho t}$ with $\rho:=f_{\nu}^{\prime}(1)=(\mathbb{E}\left[\nu\right]-1)/\mu$. The process is called subcritical, critical or supercritical according to whether $\mathbb{E}\left[\nu\right]$ is, respectively, less, equal, or larger than 1.
@@
3.2 Multitype processes
Consider now a generalisation of the process presented above which is composed of $d$ types of particles. We note that we will use bold symbols $\boldsymbol{\nu}$ for a vector, and, unless stated otherwise explicitly, $\nu_{i}$ will denote the value of its $i$th coordinate. We define the generating function in this case as
$G_{\boldsymbol{\nu}}(\boldsymbol{s}):=\sum_{\boldsymbol{n}\in\mathbb{N}_{0}^{d}}\mathbb{P}\left(\boldsymbol{\nu}=\boldsymbol{n}\right)\prod_{i=1}^{d}s_{i}^{n_{i}}\quad\text{for}\quad\boldsymbol{s}=(s_{1},\dots,s_{d}).$
The multitype branching process $\boldsymbol{Z}$ with $\boldsymbol{Z}(0)=\boldsymbol{I}^{(0)}\sim\boldsymbol{I}$, is described by:
  • At times $0=T_{0}<T_{1}<T_{2}<\dots$, iid immigrants $\boldsymbol{I}^{(0)},\boldsymbol{I}^{(1)},\boldsymbol{I}^{(2)},\dots\sim\boldsymbol{I}$ join the population described by the branching process $\boldsymbol{Z}.$
  • Each particle of type $i$ lives $Exp(1/\mu_{i})$ time after which it dies and produces new ones with distribution $\boldsymbol{\nu}_{i}$ of their counts, that is to say $({\nu}_{i})_{j}$ of type $j$.
@@ 3.2 Multitype processes
  1. 1.
    the immigrants and progeny, $I_{i}$ and $(\nu_{i})_{j}$, have finite expectation, and the latter is not a.s. constant. For some of the results we impose the following stronger restriction:
    $I_{i}\text{ and }(\nu_{i})_{j}\text{ have finite second moments.}$ (H)
  2. 2.
    the mean offspring matrix
    $M:=\bigl(\mathbb{E}\left[(\nu_{i})_{j}\right]\bigr)_{1\leq i,j\leq d}$
    is assumed primitive, i.e. there exists an integer $p\geq 1$ such that $M^{p}$ has strictly positive entries. We briefly discuss the more general decomposable (or also reducible) case in Subsection 3.3.
Similarly to the single-type case,
@@
where $\boldsymbol{Z}^{(i)}_{\times}$ are iid processes with law $\boldsymbol{Z_{\times}}$, that is, without immigration and started with $\boldsymbol{I^{(i)}}$. This gives the generating function of the process, with $G_{\boldsymbol{Z}}(t,\boldsymbol{s}):=G_{\boldsymbol{Z}(t)}(\boldsymbol{s})$ and $G_{\boldsymbol{Z}_{\times}}(t,\boldsymbol{s}):=G_{\boldsymbol{Z}_{\times}(t)}(\boldsymbol{s})$,
@@
Translating the last in the language of the Laplace functional of the random measure $\Phi$ governing the process $T_{i}$, substituting in (2.1), we get that
$G_{\boldsymbol{Z}}(t,\boldsymbol{s})=\mathcal{L}_{\Phi}(f_{t}),\quad\text{with}\quad f_{t}(x):=-\ln\left(G_{\boldsymbol{Z}_{\times}}\left(t-x,\boldsymbol{s}\right)\right)\mathbbm{1}_{\left\{x\leq t\right\}}.$ (3.2)
@@
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\mathbb{E}\left[\exp(-\langle\boldsymbol{Z}(t),\boldsymbol{s}\rangle)\right],$
where $\langle\cdot,\cdot\rangle$ is usual scalar product on $\mathbb{R}^{d}$, we would have
@@
3.3 Decomposable (Reducible) Case
In the more general decomposable (or reducible) setting, the mean offspring matrix $M$ can be permuted into a block-triangular form
$PMP^{-1}=\begin{pmatrix}M_{11}&M_{12}&\cdots&M_{1k}\\ 0&M_{22}&\cdots&M_{2k}\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&M_{kk}\end{pmatrix}$,
where each diagonal block $M_{ii}$ is irreducible (primitive). One then studies the multitype branching process in stages:
  • Block 1 (leading class): behaves as an irreducible system with its own growth rate; immigration into this class follows the point process mechanism as before.
  • Block $r$ (subsequent class): receives external immigration from class $1,\dots,r-1$ via their off-diagonal connections $M_{jr}$, in addition to the original point-process arrivals. One treats the output of earlier blocks as an inhomogeneous immigration input and applies the irreducible theory to each block in turn.
Asymptotic results for each class can then be obtained by iterating the single-block analysis, noting that the effective immigration intensity into block $r$ is a superposition of point-process-driven arrivals plus the evolving contribution from blocks $1$ through r1r-1.
@@ 4.1 Single-type processes
A standard argument using martingale theory, for example [2, Theorem 1 on p.111], shows that for the process without immigration $Z_{\times}$, there exists a real-valued random variable $W_{\times}$ such that
@@
In the supercritical case, it is natural to characterise some of the properties of $W_{\times}$, e.g., the existence of a density. In the sub- and critical cases, this $W_{\times}$ is a.s. equal to 0 and a common question is to analyse the so-called Yaglom limit as $t\to\infty$ of quantities such as $\mathbb{P}(Z_{\times}(t)>f(t)|Z_{\times}(t)>0)$ for some $f$, see [2, Chapter III, Theorems 2-4]. Similar results are also available for the case of age-dependent processes, see [2, Chapter IV].
@@
    In the case of immigration spanned by a $(\Lambda,K)$-DPP process, the expectation $\mathbb{E}\left[s^{Z(t)}\right]$ is equal to
    $1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-G_{Z_{\times}}(t-x_{i},s)\right)\Lambda(\mathrm{d}\boldsymbol{x})$ (4.3)
@@
    In the case of immigration spanned by a Cox process with a directing measure $\eta$,
    $\mathbb{E}\left[s^{Z(t)}\right]={\mathbb{E}\left[\exp\left(-\int_{(0,t]}\left(1-G_{Z_{\times}}(t-x,s)\right)\eta(\mathrm{d}x)\right)\right]}.$
@@
    In the case of immigration spanned by a $(\beta,\lambda)$-FPP,
    $\mathbb{E}\left[s^{Z(t)}\right]={\mathbb{E}\left[\exp\left(-\lambda\int_{(0,t]}\left(1-G_{Z_{\times}}\left(t-x,s)\right)\right)\mathrm{d}Y_{\beta}(x)\right)\right]},$
    where $Y_{\beta}$ is the inverse of an independent $\beta$-stable subordinator.
To obtain the respective Laplace transforms, one should replace $G_{Z_{\times}}$ with $\mathcal{L}_{Z_{\times}}$.
@@
The generating functions are directly obtained by substituting (3.2) into the relevant Laplace functional, that is, (2.4) for Cox processes, (2.6) for FPPs, and (2.3) for DPPs. For the Laplace transform, instead of (3.2), we should substitute (3.3). ∎
@@
4.2 Multitype processes
The results for multitype processes have similar nature, however generalising the one-dimensional results is not always direct and requires a careful treatment. To extend (4.1), let us define an analogue of $\mathbb{E}\left[\nu\right]$, that is $e^{A}$, where the generator $d\times d$ matrix $A$ is defined by
$A_{i,j}:=\mu_{i}^{-1}\left(\mathbb{E}\left[\left(\nu_{i}\right)_{j}\right]-\mathbbm{1}_{\left\{i=j\right\}}\right).$ (4.4)
We will suppose that $A$ is irreducible and let $\rho$ be its largest eigenvalue (also known as Perron-Frobenius root). The process is called sub-, super- or just critical according to $\rho$, respectively, less, larger or equal than 0. Let $\boldsymbol{u}$ and $\boldsymbol{v}$ be row vectors such that
$A\boldsymbol{u}^{t}=\rho\boldsymbol{u}^{t},\quad\boldsymbol{v}A=\rho\boldsymbol{v},\quad\sum_{i}u_{i}v_{i}=1,\quad\text{and}\quad\sum_{i}u_{i}=\\ 1,$
with $t$ denoting the transpose. Moreover, it is known, see [2, p.203] that the coordinates of $\boldsymbol{u}$ and $\boldsymbol{v}$ are strictly positive. Then, from [2, Theorem 2, p.206], rewritten in the form [36, Lemma 2.1], there exists a real-valued random variable $W_{\times}$ such that
@@ 4.2 Multitype processes
$G_{\boldsymbol{Z}_{\times}}(\boldsymbol{s})=G_{\boldsymbol{I}}\left(G_{\boldsymbol{Z}^{(1)}_{\times}}(\boldsymbol{s}),\dots,G_{\boldsymbol{Z}^{(d)}_{\times}}(\boldsymbol{s})\right),$
@@
The statements of Theorem 4.1 extend to the vector case by $s\to\boldsymbol{s}$, $Z\to\boldsymbol{Z}$, in all regimes (subcritical, critical, supercritical). The same holds if one replaces generating functions by Laplace transforms $G_{\boldsymbol{Z}}\to\mathcal{L}_{\boldsymbol{Z}}$ and $G_{\boldsymbol{Z}_{\times}}\to\mathcal{L}_{\boldsymbol{Z}_{\times}}$.
@@
As noted previously, the case of immigration generated by a Poisson random measure is analysed ’in [17, 44, 45] and independently by Rabehasaina and Woo in [36]. For immigration times that form a renewal process, see [43]. The latter may be applied in the case of an immigration spanned by a FPP after checking that the used arguments can be modified to include the case of an infinite expectation of the inter-arrival times.
As an illustrative example of how ideas from the case of Poisson immigration can be modified to encompass the case of immigration spanned by a DPP, we generalise [36, Theorems 3, 4, and 5] (since the Poisson point process is a DPP), which are also available in [17, 44, 45].
5 Multitype branching processes with DPP immigration
@@ 5.1 Moments of the process
Let us recall that for a vector $\boldsymbol{Z}$, we denote by $Z_{i}$ its $i$th coordinate. Then we have the following.
@@
In the case of immigration spanned by a $(\Lambda,K)$-DPP,
  1. 1.
@@
This follows by differentiating the generating function of the process, derived as the multidimensional version of (4.3). We provide the complete calculation in Section 6.1.
5.2 Asymptotics
@@
Let $\rho$ be the Perron-Frobenius root of the matrix $A$. In the case of immigration spanned by a $(\Lambda,K)$-DPP $\Phi$ with kernel $K$ such that $\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)$ is finite, there exists an $\mathbb{R}^{d}$-valued random variable $\boldsymbol{W}$ such that
$\frac{\boldsymbol{Z}(t)}{e^{\rho t}}\xrightarrow[t\to\infty]{d}\boldsymbol{W}.$ (5.1)
@@
$\boldsymbol{W}\stackrel{{\scriptstyle d}}{{=}}\sum_{i}\boldsymbol{v}\cdot W^{(i)}_{\times}e^{-\rho T_{i}},\quad\text{and}\quad\mathbb{E}\left[\exp\left(-\langle\boldsymbol{W},\boldsymbol{s}\rangle\right)\right]=\mathcal{L}_{\Phi}\left(-\ln\left(\mathcal{L}_{\boldsymbol{v}W_{\times}}\left(\boldsymbol{s}e^{-\rho x}\right)\right)\right),$
where $W^{(i)}_{\times}$ are iid copies of $W_{\times}$ and $T_{i}$ being the atoms of $\Phi$.
@@
where $\boldsymbol{x}+t:=(x_{1}+t,\dots,x_{n}+t)$; see [3, Chapter 5.6]. This property holds, for example, if the kernel is shift-invariant, K(x,y)=K(x+t,y+t)=:C(xy)K(x,y)=K(x+t,y+t)=:C(x-y), as in the case of the Poisson kernel, $K(x,y)=\mathbbm{1}_{\left\{x=y\right\}}$, or the Ginibre (Gaussian) kernel: $K(x,y)=\exp\left(-(x-y)^{2}/2\right)/\pi$. A result by Lavancier et al. [24, Theorem 1] states that $C$ can be any continuous correlation function with eigenvalues in its spectral representation bounded by 1, with examples provided in their work.
Before stating our next result, we introduce the asymptotic equivalence notation, provided $g(t)$ is non-zero for sufficiently large $t$,
$f(t)\stackrel{{\scriptstyle\infty}}{{\sim}}g(t)\quad\text{if}\quad\lim_{t\to\infty}\frac{f(t)}{g(t)}=1.$
@@
Assume that immigration is spanned by a stationary $(\Lambda,K)$-DPP with $\Lambda(\mathrm{d}x)=\lambda(x)\mathrm{d}x$, and that $\lambda(x)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}e^{\delta x}$ for some $\lambda_{\infty}>0$ and $\delta\in\mathbb{R}$.
@@
    Under (H), if $\delta>\max\{\rho,0\}$, then
    $\frac{{Z_{i}}(t)}{e^{\delta t}}\xrightarrow[t\to\infty]{L^{2}}A_{i}:=K_{\ast}\lambda_{\infty}\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x,$
@@
    Under (H), if the process is supercritical, that is, $\rho>0$, and $\delta=\rho$, then
@@
    $\frac{\boldsymbol{Z}(t)}{t}\xrightarrow[t\to\infty]{d}Y\boldsymbol{v},$
    where $Y\sim\Gamma(K(0,0)\lambda_{\infty}\beta,1/Q)$ with
    $Q:=\frac{1}{2}\sum_{i,j,k=1}^{d}\left.\frac{\partial^{2}G_{\nu_{i}}}{\partial x_{j}\partial x_{k}}\right|_{\boldsymbol{x}=\boldsymbol{1}}\mu_{i}^{-1}v_{i}u_{j}u_{k},\qquad\beta:=\frac{\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle}{Q},$
@@
The result of item 3 in the last theorem in the case of Poisson immigration is established in [17, Theorem 8] and, under the stronger assumption that all moments of $I_{i}$ exist, in [36, Theorem 3]. After a close inspection, we have observed that the two in fact do not match. This is due to two errata: in the proof of [17, Theorem 8], there is a missed $r^{\ast}$ on the second line of page 221. The error in [36, Theorem 3] is due to an incorrect application to the result of Weiner [49]. In his work, the eigenvectors $\boldsymbol{u}_{W}$ and $\boldsymbol{v}_{W}$ are defined as, respectively, the right and left eigenvector of the matrix composed of $\mathbb{E}\left[(\nu_{i})_{j}\right]-\mathbbm{1}_{\left\{i=j\right\}}$, and not as the eigenvectors of the generator matrix $A$, see (4.4). The normalisation is $\langle\boldsymbol{u}_{W},\boldsymbol{1}\rangle=1$ and $\langle\boldsymbol{u}_{W},\boldsymbol{v}_{W}\rangle=1$. It is then direct that,
$u_{W,i}:=(u_{W})_{i}=u_{i},\quad\!v_{W,i}=\frac{1}{\sum_{j}(u_{j}v_{j}/\mu_{j})}\frac{v_{i}}{\mu_{i}},\quad\!\text{and}\quad\!\sum_{i}\mu_{i}u_{W,i}v_{W,i}=\frac{1}{\sum_{i}(u_{i}v_{i}/\mu_{i})}.$ (5.3)
@@
The theorem is proved in Section 6.3. For the proof of the first two items, we will need the following estimates.
Lemma 5.5.
Under the assumptions of Theorem 5.3 and (H),
  1. 1.
    if $\delta>\rho$, then for some $A_{i}\in\mathbb{R}_{+}$,
    $\mathbb{E}\left[Z_{i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}A_{i}e^{\delta t},\quad\text{and}\quad\operatorname{Var}(Z_{i}(t))=\begin{cases}\mathrm{O}\left(e^{\max\{2\rho,\delta\}t}\right)&\text{if $\rho>0$ and $\delta\neq 2\rho$},\\ \mathrm{O}\left(te^{\delta t}\right)&\text{if $2\rho=\delta>0$}.\end{cases}$
  2. 2.
    if $\delta=\rho>0$, then for some $A_{i}^{\prime}\in\mathbb{R}_{+}$,
    $\mathbb{E}\left[Z_{i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}A^{\prime}_{i}te^{\delta t},\quad\text{and}\quad\operatorname{Var}(Z_{i}(t))=\mathrm{O}\left(e^{2\delta t}\right).$